### v1

@@ 5 Constraint Types
Let the two districts be $L$ and $R$. We start by lower bounding the expected number of ordered pairs of students $(i_{1},i_{2})$ that satisfy the following conditions:
  1. 1.
  2. 2.
    $i_{1}$ most prefers a school $\ell\in L$.
  3. 3.
    $i_{2}$ most prefers a school $r\in R$, and next prefers $\ell$.
  4. 4.
    No other student finds $\ell$ or $r$ acceptable.
  5. 5.
    $i_{2}$ is higher in the priority order than $i_{1}$ at $\ell$.
Under these conditions, $i_{1}$ would strictly benefit from $i_{2}$ becoming unconstrained, as $i_{2}$ would switch from enrolling in district $L$ to enrolling in district $R$, which would open up school $\ell$ for $i_{1}$.
We want to determine the probability that this set of conditions occurs for a specific $i_{1}$ and $i_{2}$. Conditioned on all previous conditions, the first condition occurs with probability at least $p^{2}$, the second condition occurs with probability at least $\tfrac{1}{2}$, and the third condition occurs with probability $\tfrac{n-1}{n}\cdot\tfrac{1}{2}\cdot\tfrac{1}{n-1}$. The fourth condition occurs with probability $\left(\frac{n-2}{n}\cdot\frac{n-3}{n-1}\right)^{n-2}$. Finally, the fifth condition occurs with probability $\tfrac{1}{2}$. For sufficiently large $n$, this probability can be lower bounded by
$\frac{p^{2}}{2^{3}}\cdot\frac{1}{n}\cdot\left(\frac{n-2}{n}\cdot\frac{n-3}{n-1}\right)^{n-2}\geq\frac{p^{2}}{2^{4}e^{4}(n-1)}.$
The number of ordered pairs of students is $n(n-1)$. List all such pairs and let the $j$th element of the list be $I_{j}$. Let $\mathbbm{1}_{j}$ be the indicator random variable which represents whether the students in $I_{j}$ satisfy all of the above conditions. Then the expected number of pairs of students that satisfy all conditions is
To conclude the proof, we observe that by construction, no student can participate in two different pairs that satisfy the above conditions. Therefore, all such pairs of students involve distinct students, so there are at least an expected $\frac{p^{2}}{2^{4}e^{4}}\cdot n$ unconstrained students who each strictly prefer for a distinct constrained student to become unconstrained. We choose $\tau=\frac{p^{2}}{2^{4}e^{4}}$ to satisfy the theorem statement. ∎
We now demonstrate the use of the second, slightly more elaborate, construction.
Lemma 5.3.
@@
For every $p\in(0,1)$, there exists $\tau>0$ such that for any large enough $n$, in the uniform $(n;2)$ model with the districts using any combination of matching mechanisms, there exists a set of constrained students of expected size at least $\tau n$ where each constrained student strictly prefers for a distinct constrained student in another district to become unconstrained.
Proof.
Let the two districts be $L$ and $R$. We start by lower bounding the expected number of ordered triplets of students $(i_{1},i_{2},i_{3})$ that satisfy the following conditions:
  1. 1.
  2. 2.
    $i_{1}$ most prefers a school $\ell_{1}\in L$.
  3. 3.
    $i_{2}$ most prefers a school $\ell_{2}\in L$, and next most prefers a school $r_{1}\in R$.
  4. 4.
    $i_{3}$ most prefers $r_{1}$, and next most prefers $\ell_{1}$.
  5. 5.
    No other student finds $\ell_{1},\ell_{2}$, or $r_{1}$ acceptable.
  6. 6.
Under these conditions, $i_{1}$ would strictly benefit from $i_{2}$ becoming unconstrained, as $i_{2}$ would switch from enrolling in district $R$ to enrolling in district $L$. As a result, $i_{3}$ would switch from enrolling in district $L$ to enrolling in district $R$, which would open up school $\ell_{1}$ for $i_{1}$.
We want to determine the probability that this set of conditions occurs for a specific $(i_{1},i_{2},i_{3})$. Conditioned on all previous conditions, the first condition occurs with probability at least $p^{3}$, the second condition occurs with probability at least $\tfrac{1}{2}$, the third condition occurs with probability $(\tfrac{n-1}{n}\cdot\tfrac{1}{2}\cdot\tfrac{n-2}{n-1}\cdot\tfrac{1}{2})$, and the fourth condition occurs with probability $(\tfrac{1}{n}\cdot\tfrac{1}{n-1})$. The fifth condition occurs with probability $\left(\frac{n-3}{n}\cdot\frac{n-4}{n-1}\right)^{n-3}$. Finally, the sixth condition occurs with probability $\tfrac{1}{2^{2}}$. For sufficiently large $n$, this probability can be lower bounded by
$\frac{p^{3}}{2^{5}}\cdot\left(\frac{n-1}{n}\cdot\frac{n-2}{n-1}\right)\cdot\left(\frac{1}{n}\cdot\frac{1}{n-1}\right)\cdot\left(\frac{n-3}{n}\cdot\frac{n-4}{n-1}\right)^{n-3}\geq\frac{p^{3}}{2^{6}e^{6}(n-1)(n-2)}.$
The number of ordered triplets of students is $n(n-1)(n-2)$. List all such triplets and let the $j$th element of the list be $I_{j}$. Let $\mathbbm{1}_{j}$ be the indicator random variable that represents whether the students in $I_{j}$ satisfy all of the above conditions. Then the expected number of triplets of students that satisfy all conditions is
To conclude the proof, we observe that by construction, no student can participate in two different triplets that satisfy the above conditions. Therefore, all such triplets of students involve distinct students, so there are at least an expected $\frac{p^{3}}{2^{6}e^{6}}\cdot n$ constrained students who each strictly prefer for a distinct constrained student in another district to become unconstrained. We choose $\tau=\frac{p^{3}}{2^{6}e^{6}}$ to satisfy the theorem statement. ∎
@@ 6 Discussion
In this paper, we show that several key results regarding sincere and sophisticated students from the seminal paper of Pathak and Sönmez (2008) no longer hold in a multi-district setting. This highlights the importance of weighing the specifics of the market in question when designing centralized mechanisms, where “the market in question” should be defined very broadly, perhaps more so than customarily considered.
Several aspects of our model and results would potentially benefit from further research. First, although the idea of district choice is motivated by students and their families boundary hopping or paying a premium to move, we do not explicitly model the associated (financial or risk-taking) costs. Future research could impose a price on choosing a district other than one’s own district of residence, which would then factor into students’ strategic considerations. For this to be effective, it would also be necessary to think about student satisfaction with different school assignments using cardinal utilities rather than ordinal preferences; modeling such costs is therefore outside the scope of this paper.

### v2

@@ 6.1.1 Example: When all sincere students are constrained
Let there be two districts with schools $\ell_{1},\ell_{2},\ell_{3}\in L$ and schools $r_{1},r_{2},r_{3},r_{4}\in R$, where each school has unit capacity. District $L$ uses BM, while district $R$ uses DA. Furthermore, let there be six students $i_{1}$, $i_{2}$, $i_{3}$, $i_{4}$, $i_{5}$, and $i_{6}$.
The students’ sophistication types and preference orderings, as well as the schools’ priority orderings, are shown in the table below. Students whose preference orderings contain schools in only one district reside in that district and are constrained to it. Other students (those whose preference orderings contains schools in both districts) are unconstrained and reside in an arbitrary district.
Student preferences
$\text{(sincere) }i_{1}:$ $\ell_{1}\succ\ell_{2}$
$\text{(sophisticated) }i_{2}:$ $\ell_{2}\succ r_{3}$
$\text{(sophisticated) }i_{3}:$ $r_{1}\succ r_{2}$
$\text{(sophisticated) }i_{4}:$ $r_{2}\succ r_{1}$
$\text{(sophisticated) }i_{5}:$ $r_{1}\succ r_{3}\succ\ell_{3}$
$\text{(sincere) }i_{6}:$ $\ell_{1}$
School priority orderings
$\ell_{1}:$ $i_{6}-i_{1}$
$\ell_{2}:$ $i_{1}-i_{2}$
$\ell_{3}:$ $i_{5}$
$r_{1}:$ $i_{4}-i_{5}-i_{3}$
$r_{2}:$ $i_{3}-i_{4}$
$r_{3}:$ $i_{2}-i_{5}$
We will show that $i_{3}$ and $i_{4}$ prefer for $i_{1}$ to become sophisticated. First, consider the original setting where $i_{1}$ is sincere (and constrained). Since $i_{6}$ both ranks $\ell_{1}$ highest and has the highest priority at that school among all students who find it acceptable, $i_{6}$ is assigned to $\ell_{1}$. Observe that the sophisticated-unconstrained student $i_{2}$ enrolls in district $L$, since she is thus assigned to $\ell_{2}$ as the only student who applies to this school in the first round. Therefore, $i_{1}$ remains unassigned. Sophisticated-unconstrained student $i_{5}$ enrolls in district $R$ because this gets her assigned to school $r_{3}$, which she strictly prefers to any district $L$ school. The matching process results in $i_{1}$ unassigned, $i_{2}$ assigned to $\ell_{2}$, $i_{3}$ assigned to $r_{2}$, $i_{4}$ assigned to $r_{1}$, $i_{5}$ assigned to $r_{3}$, and $i_{6}$ assigned to $\ell_{1}$.
Suppose instead that $i_{1}$ becomes sophisticated. By the same argument, $i_{6}$ is assigned to $\ell_{1}$. Diverging from the previous analysis, $i_{1}$ applies to school $\ell_{2}$ in the first round and gets assigned to it. $i_{2}$ therefore enrolls in district $R$ rather than $L$, and since she has the highest priority at $r_{3}$ among all students who find it acceptable, $i_{2}$ is assigned to $r_{3}$. Therefore, since enrolling in district $R$ would now leave student $i_{5}$ unassigned, she enrolls in district $L$ and is assigned to $\ell_{3}$. The matching process results in $i_{1}$ assigned to $\ell_{2}$, $i_{2}$ assigned to $r_{3}$, $i_{3}$ assigned to $r_{1}$, $i_{4}$ assigned to $r_{2}$, $i_{5}$ assigned to $\ell_{3}$, and $i_{6}$ assigned to $\ell_{1}$. Therefore, both $i_{3}$ and $i_{4}$ strictly prefer for $i_{1}$ to become sophisticated. ∎
6.1.2 Shortcomings of Large-Market Analysis without Sincere-Unconstrained Students
We observe that replacing the gadget in our current proof of Theorem 3.1 by the school choice problem from Proposition 6.1 results in a weaker guarantee. Specifically, the resulting guarantee is of an expected number of sophisticated students who prefer for some sincere student to become sophisticated that is only constant, rather than a constant fraction of the total number of students. [^17] This is due to the existence of a cycle in the preferences of $i_{4}$ and $i_{5}$, similar to the one in the preferences of $i_{1}$ and $i_{2}$ in the proof of Theorem 4.1. We thus get a quantitative guarantee similar to that of Theorem 4.1 rather than to that of Theorem 3.1.
Consider a school choice problem in which some sophisticated student strictly prefers for a sincere student to become sophisticated. We call this school choice problem weak if using it to replace the gadget in the proof of Theorem 3.1 results in a constant-number or weaker guarantee. We call this school choice problem strong if using it to replace the gadget in the proof of that theorem results in the constant-fraction guarantee as in the statement of that theorem. In particular, the school choice problem from Section 3.1 is strong and the school choice problem from the proof of Proposition 6.1 is weak.
We now argue that being weak is a limitation not only of the school choice problem from the proof of Proposition 6.1, but rather of every school choice problem that lacks sincere-unconstrained students. We start by investigating which aspects of the machinery of Pathak and Sönmez (2008) transfers to our setting, in what way, and with what limitations.
Fix some multi-district school choice problem (possibly containing sincere-unconstrained students). Without loss of generality, assume that each constrained student only has in its preference list schools that are in its district. Generalizing a construction by Pathak and Sönmez (2008), define the following single-district school choice problem, to which we refer as the “augmented economy.” The schools and students in the augmented economy are identical to those in the original, multi-district school choice problem, but the priorities of each school originally located in a district that uses BM is modified as follows to partition students into tiers, where within each tier the priorities are the same as in the original school choice problem, but each student in the first tier has priority over each student in the second tier, and so on. For a school $s$, the first tier consists of all sophisticated students, as well as all sincere students who prefer $s$ over all other schools in the same district as $s$ in the original multi-district school choice problem. [^18] Recall that students who are originally constrained to districts other than the school’s do not have this school in the preference list, so while such sophisticated students are included in this tier, this is inconsequential. For every $j\geq 2$, the $j$th tier consists of all sincere students who strictly prefer precisely j1j-1 schools from the original district of $s$ over $s$.
The proof of Proposition 4 of Pathak and Sönmez (2008) (which precludes the possibility of a result like our Theorems 3.1 and 6.1 in their model) hinges on two properties of the augmented economy in that paper: First, the set of equilibrium outcomes of the original school choice problem coincides with the set of stable matchings of the augmented economy (this is their Proposition 1); and second, every school that participates in BM gives priority in the augmented economy to all sophisticated students, as well as all sincere students who prefer the school most, over all other students (this is a condition of their Lemma 1). The latter property is violated only by sincere-unconstrained students in our augmented economy for multi-district school choice problems. [^19] Indeed, this property is precluded by the trait discussed in the beginning of Section 6.1.1 as driving all of the constructions from Section 3: that a sincere-unconstrained student can block a sophisticated student at some school $s$ (and hence have top-tier priority at $s$), and yet, this sincere-unconstrained student may vacate that seat for another school if she becomes sophisticated (and hence, when sincere, does not have top-tier priority at a school she prefers over $s$). When all sincere students are constrained, this property is satisfied; however, the first property might be violated, as we now discuss.
An equilibrium in our original multi-district school choice problem need not be stable in the augmented economy: Such an equilibrium might contain blocking pairs with respect to the augmented economy, [^20] A blocking pair is a student–school pair where the student prefers the school to her current assignment, and the student has priority (in our case, in the augmented economy) at the school over another student currently assigned to the school. but only with a very specific structure, to which we refer as incentive-compatible (IC) blocking pairs. An IC blocking pair in an equilibrium of our original multi-district school choice problem is a student–school pair that is a blocking pair with respect to the augmented economy; and involves a school $s$ in a district that uses DA, and a (sophisticated) unconstrained student $i$ who enrolls (in the equilibrium) in a different district from $d(s)$, such that if $i$ were to deviate from the equilibrium and enroll in $d(s)$, then in the run of DA in $d(s)$, after $i$ applies to $s$ (knocking out the equilibrium match of $s$), a rejection cycle is initiated that eventually causes $i$ to be rejected from $s$. An example of an IC blocking pair can be seen in the school choice problem in the proof of Proposition 6.1 when $i_{1}$ is sophisticated: $i_{5}$ (who in equilibrium enrolls in district $L$) blocks (in the augmented economy) with school $r_{1}$ (which is in district $R$, which uses DA). However, if $i_{5}$ were to enroll in district $R$, then in the run of DA in that district, after $i_{5}$ knocks out $i_{3}$ from $r_{1}$, a rejection cycle is initiated that eventually causes the rejection of $i_{5}$ from $r_{1}$ (in favor of $i_{4}$).
In the absence of both sincere-unconstrained students and IC blocking pairs, the proof of Pathak and Sönmez (2008) in fact goes through in our setting as well. In a precise sense, therefore, sincere unconstrained students and IC blocking pairs are the only two phenomena that might preclude the result of Pathak and Sönmez (2008) in our setting; in Section 3 we demonstrated how the former might do so, and in Proposition 6.1 we demonstrated this for the latter. Therefore, any school choice problem that proves Proposition 6.1 despite not containing sincere-unconstrained students must satisfy that its equilibrium matching exhibits an IC blocking pair. [^21] Furthermore, such a school choice problem must have, in addition to a district that uses DA as mandated by the definition of an IC blocking pair, a district that uses BM for the sincere-turned-sophisticated student to reside in, because if this student resides in a district that uses DA, her behavior is unchanged when becoming sophisticated. In particular, the preferences of the students must allow for a rejection cycle, which in turn means that there must exist a cycle of students such that for every two consecutive students in the cycle, there must exist a school that both students have in their preference orderings. We now argue that such a cycle of students necessarily implies the school choice problem is weak.
Fix some multi-district school choice problem that satisfies Proposition 6.1. Consider a hypergraph $H=(V,E)$ where $V$ are the students in the school choice problem and the hyperedges are in one-to-one correspondence to schools, where the hyperedge corresponding to a specific school contains all of the vertices (students) that have this school in their preference lists. Note that students in distinct connected components of $H$ cannot affect each other, so it’s enough to consider each connected component that satisfies Proposition 6.1 separately and prove that its induced sub-problem is weak (connected components without sophisticated students who strictly prefer for some sincere students to become sophisticated can be ignored). We therefore assume henceforth that $H$ is connected.
In the probabilistic analysis in the proof of Theorem 3.1 (as well as in the proof of our other probabilistic results), the order of magnitude of our bound on the expected number of occurrences of the gadget is completely determined by the corresponding hypergraph $H$: Starting with a constant, for every vertex $v\in V$ our bound on the expected number of occurrences gains a factor of $n$, and for every hyperedge $e\in E$ this bound loses a factor of $n^{|e|-1}$. Therefore, if $|V|\leq\sum_{e\in E}\bigl(|e|-1\bigr)$, then an expected number of occurrences that is a constant fraction of $n$ cannot be guaranteed using this proof, implying that the school choice problem that we fixed is weak. Note that the cycle of students that is guaranteed above to exist is in fact a cycle in $H$. As such, $H$ satisfies $\sum_{e\in E}\bigl(|e|-1\bigr)\geq|V|$ (recall that for every connected hypergraph, $\sum_{e\in E}\bigl(|e|-1\bigr)\geq|V|-1$, with equality only for hypertrees, i.e., for acyclic connected hypergraphs). Therefore, the school choice problem that we fixed is weak.
Overall, we have shown that no school choice problem in which all sincere students are constrained can be used as the gadget in our proof of Theorem 3.1. While this technically does not rule out other proof techniques, it is a strong indication of a key role that sincere-unconstrained students play in the phenomenon formulated by Theorem 3.1. Given this result, in the next section we ask how robust Theorem 3.1 is to the precise definition of sincere-unconstrained students.
6.2 Robustness to the Definition of Sincere-Unconstrained Students
The definition of sincere-unconstrained students is not beyond criticism. Indeed, one might wonder why a student who possesses sufficient understanding of the intricacies of the mechanisms used by the different districts (and the preferences of other students) so as to be able to perfectly optimize between them does not realize the potential benefits of reporting strategically within the district in which she chooses to enroll. Having ascertained the key role of sincere-unconstrained students in the proof of Theorem 3.1, it is therefore natural to ask how robust Theorem 3.1 is to weakening the ability of such students to fully optimize between districts. Specifically, sincere-unconstrained students might be more realistically modeled as students who are able (e.g., financially) to enroll in a district of their choice, but do not fully grasp the details of the mechanism used in each district and the preferences of other students (and hence report truthfully within the district that they end up choosing). Such students, therefore, are unable to fully optimize their choice of district and have to resort to using some heuristic to choose between districts. In this section, we consider sincere-unconstrained students who choose between districts based on a natural broad class of heuristics that we term positional heuristics. We show that regardless of the specific heuristics within this broad class that are used, the main results of the previous sections still hold.
There are many reasonable heuristics that sincere-unconstrained students might employ to choose between districts, and different heuristics may lead to different choices of districts to enroll in. For example, a sincere-unconstrained student could always enroll in the district of her first-choice school (and then report her true preferences over schools in that district). This can, of course, be suboptimal: Such a sincere-unconstrained student who can never get assigned to her first-choice school will still enroll in the district of that school; in contrast, a sincere-unconstrained student who fully strategizes over her district choice (like those considered in the previous section) would instead enroll in another district if she would get a preferable school assignment by reporting her true preferences over schools in that district.
Beyond fully strategizing over district choice and always enrolling in the district of her first-choice school, there are many other reasonable ways a sincere-unconstrained student can choose where to enroll. For example, a sincere-unconstrained student might simply count the number of schools on her preference ordering from each district, and enroll in the district with the most such schools. Interpolating between this and considering only her first-choice school, she could enroll in the district that has the most schools in, e.g., the top ten schools of her preference ordering. Alternatively, she might score each school based on how high up it is in her preference ordering, e.g., by giving her least preferred (but still acceptable) school one point, her next least preferred school two points, and so on, and then enrolling in the district with the highest sum of points over all schools.
We believe that all of these definitions have merit in different contexts, and therefore avoid choosing between them. Instead, we show a version of our results that holds for a large class of heuristics for sincere-unconstrained students that include all of the above and many more. Taking inspiration from social choice theory (e.g., Young, 1975; Boutilier et al., 2012), we define the class of scoring functions as follows.
Definition 6.2.
  • For $v\in\mathbb{R}^{n}$, define the scoring function $f^{v}$ as follows. For any preference ordering $P_{i}$, define $P_{i}(s)$ as the ranking of school $s$ in $P_{i}$. Further let $v\bigl(P_{i}(s)\bigr)=v_{P_{i}(s)}$ if $s\in P_{i}$ and $v\bigl(P_{i}(s)\bigr)=0$ otherwise. Then we define the scoring function as
[…]
  • A scoring function is monotone if both $v\bigl(P_{i}(s)\bigr)\geq v\bigl(P_{i}(s^{\prime})\bigr)$ whenever $s$ is ranked higher than $s^{\prime}$ in $P_{i}$ and $v\bigl(P_{i}(s)\bigr)\geq 0$ whenever $s\in P_{i}$.
  • A sincere-unconstrained student uses a positional heuristic if she chooses in which district to enroll based on a monotone scoring function. In the case of a tie between districts, we assume that a student using a positional heuristic enrolls in the district of the school that she ranks first among all schools in the tied districts.
We emphasize that all of the heuristics discussed above are positional heuristics. For example, one positional heuristic that corresponds to a sincere-unconstrained student choosing to enroll in the district of her first-choice school is the heuristic that assigns a score of $1$ to the first-choice school and a score of 0 to every other school. If a sincere-unconstrained student instead enrolls in the district with the most schools they find acceptable, this is equivalent to using a positional heuristic that assigns a score of $1$ to every school in her preference ordering (and 0 to every other school). If a student scores each school by how far it is from the end of her preference ordering, this is equivalent to using the positional heuristic that assigns a score of [length of ROL+1][$\text${length of ROL}+1-$\ell$] to the school that is ranked $\ell$ on her list and is analogous to the well-known Borda Count voting rule (Black, 1958).
As we show in this section, our main results also hold when each sincere-unconstrained student uses some positional heuristic to choose between districts (with different sincere-unconstrained students possibly using different heuristics or even fully optimizing). We note that our results in Sections 4 and 5 do not require sincere-unconstrained students at all. Therefore, the results in those sections hold regardless of how a sincere-unconstrained student chooses between districts (even if a sincere-unconstrained student does not use a monotone scoring function). We therefore need only justify how our example and theorem from Section 3 continue to hold when sincere-unconstrained students use positional heuristics.
We first adapt the example from Section 3.1. Recall that there is only one sincere-unconstrained student in this example, whose preferences are $r_{2}\succ r_{1}\succ\ell_{1}$, and who, when fully optimizing between districts but reporting truthfully within a district, enrolls in district $L$ (and causes $i_{2}$ to not get assigned to school $\ell_{1}$). Our key insight is that by adding two schools in district $L$ to the top of the preference list of this student, we can guarantee that she will always choose to enroll in district $L$, regardless of the monotone scoring function used (and consistent with full optimization). This is because by monotonicity, the score of each school in district $R$ is at most the score of the school ranked two above it in district $L$, so the overall score for district $R$ must be at most the score for district $L$. In order to ensure that $i_{1}$ is not assigned to these new district-$L$ schools when her type is sincere-unconstrained, we also need to add two students who reside in district $L$, each having one of the new schools as her top choice (and each having top priority at that respective school). The outcomes of this modified example match the outcome in Example 3.1, except that the two new students are also assigned to the two new schools. To ensure that $i_{1}$, when enrolling in district $L$, still causes $i_{2}$ to not get assigned to school $\ell_{1}$, we have district $L$ use DA. The full example is given in Appendix C.
Finally, we adapt our large-market analysis to show that a version of Theorem 3.1 still holds when sincere-unconstrained students use positional heuristics, which we prove in Appendix C.
Theorem 6.3.
@@ 6.2 Robustness to the Definition of Sincere-Unconstrained Students
Suppose that every sincere-unconstrained student uses a positional heuristic to decide which district to enroll in. Then for every $p\in(0,1)$, there exists $\tau>0$ such that for any large enough $n$, in the uniform $(n;5)$ model with category-probability lower bound $p$ and with one district using BM and the other using DA, there exists a set of sophisticated students of expected size at least $\tau n$ where each sophisticated student strictly prefers for a distinct sincere student to become sophisticated, and weakly prefers for all other sincere students to become sophisticated.
@@ 7 Discussion
In this paper, we show that several key results regarding sincere and sophisticated students from the seminal paper of Pathak and Sönmez (2008) no longer hold in a multi-district setting. This highlights that when weighing the specifics of the market in question when designing centralized mechanisms, it might be useful to define “the market in question” more broadly than is customary.
Several aspects of our model and results would potentially benefit from further research. First, although the idea of district choice is motivated by students and their families boundary hopping or paying a premium to move, we do not explicitly model the associated (financial or risk-taking) costs. Future research could impose a price on choosing a district other than one’s own district of residence, which would then factor into students’ strategic considerations. For this to be effective, it would also be necessary to think about student satisfaction with different school assignments using cardinal utilities rather than ordinal preferences; modeling such costs is therefore outside the scope of this paper. [^22] A recent paper by Artemov and Tomoeda (2025) studies a single district in which students can have “walking distance” priority for a school based on where they reside. Future research into multi-district school choice might draw inspiration from the way the model of Artemov and Tomoeda incorporates a moving cost.